### v2

# MomentSeeker: A Comprehensive Benchmark and A Strong Baseline For Moment Retrieval Within Long Videos

Abstract
Retrieval augmented generation (RAG) holds great promise in addressing challenges associated with long video understanding. These methods retrieve useful moments from long videos for their presented tasks, thereby enabling multimodal large language models (MLLMs) to generate high-quality answers in a cost-effective way. In this work, we present MomentSeeker, a comprehensive benchmark to evaluate retrieval models’ performance in handling general long-video moment retrieval (LVMR) tasks. MomentSeeker offers three key advantages. First, it incorporates long videos of over 500 seconds on average, making it the first benchmark specialized for long-video moment retrieval. Second, it covers a wide range of task categories (including Moment Search, Caption Alignment, Image-conditioned Moment Search, and Video-conditioned Moment Search) and diverse application scenarios (e.g., sports, movies, cartoons, and ego), making it a comprehensive tool for assessing retrieval models’ general LVMR performance. Additionally, the evaluation tasks are carefully curated through human annotation, ensuring the reliability of assessment. We further fine-tune an MLLM-based LVMR retriever on synthetic data, which demonstrates strong performance on our benchmark. We perform extensive experiments with various popular multimodal retrievers based on our benchmark, whose results highlight the challenges of LVMR and limitations for existing methods. Our created resources is available at https://github.com/yhy-2000/MomentSeeker.
MomentSeeker: A Comprehensive Benchmark and A Strong Baseline For Moment Retrieval Within Long Videos
Huaying Yuan1, Jian Ni3, Yueze Wang1, Junjie Zhou4, Zhengyang Liang2, Zheng Liu2*, Zhao Cao3, Zhicheng Dou1*  and Ji-Rong Wen1* 1Gaoling School of Artificial Intelligence, Renmin University of China 2Beijing Academy of Artificial Intelligence 3School of Smart Governance, Renmin University of China 4Beijing University of Posts and Telecommunications {hyyuan, dou}@ruc.edu.cn
1 Introduction
Figure 1: Demonstration of our benchmark. Given a natural language query $\textbf{q}_{\textbf{T}}$ (optionally accompanied by an image $\textbf{q}_{\textbf{I}}$ or video $\textbf{q}_{\textbf{V}}$), models should search the provided video to precisely identify the most relevant clips. We marked the parts that directly answer the query with red circles.
Long video understanding remains a significant challenge Fu et al. (2024); Zhou et al. (2025); Wang et al. (2024b). Retrieval augmented generation (RAG) Luo et al. (2024); Ataallah et al. (2024) holds great promise in addressing challenges associated with long video understanding. These methods retrieve useful moments from long videos for their presented tasks, thereby enabling multi-modal large language models (MLLMs) to generate high-quality answers in a cost-effective way. A key research problem is how to accurately retrieve essential information to answer a given query based on user instructions. In real word search scenarios such as video agent Fan et al. (2024); Wang et al. (2024c), users may ask “Where did I put my phone?”, which requires the model to analyze the complete historical context to pinpoint relevant moments. These emerging demands underscore the growing importance of video retrieval in practical applications. In this paper, we formally define the task of precisely locating key information segments within a long video based on a given query as long-video moment retrieval (LVMR).
Existing video retrieval benchmarks Xu et al. (2016); Chen and Dolan (2011); Lei et al. (2021); Hendricks et al. (2017) evaluate cross-modal retrieval by using video captions as text queries to retrieve relevant videos from a candidate pool. However, these benchmarks are primarily designed for short videos, resulting in a significant disparity from real-world long-video search scenarios. Practical applications often require capabilities such as retrieving discontinuous events within lengthy context, which remains unaddressed in current frameworks. Furthermore, these benchmarks rely on sentence-like captions as queries, which presents a marked contrast with real-world retrieval scenarios where queries are typically user-generated (e.g., natural language questions).
To address these gaps, we present MomentSeeker, a comprehensive benchmark specifically designed for evaluating long-video moment retrieval (LVMR) capabilities. It introduces three fundamental innovations:
(1) First LVMR-Specialized Benchmark. Existing multi-modal retrieval benchmarks predominantly focus on short video clips, failing to address the unique challenges of LVMR where systems must localize precise moments within hour-long videos. MomentSeeker pioneers long-video understanding with an average video duration of over 500 seconds and a maximum of 7100 seconds, requiring the ability to distinguish subtle temporal relationships within continuous footage. This contrasts with conventional cross-video retrieval paradigms, simulating practical applications like retrieval-augmented video understanding.
(2) Comprehensive Task Coverage. As shown in Figure 1, our benchmark spans 4 meta-tasks and 18 sub-tasks across three dimensions: 1) Question types covering both descriptive captions and naturalized instructional queries (e.g. “Where was the weighing scale?”), which is more practical in real-world search scenarios; 2) Multi-modal Query Fusion, supporting text, image, and video conditionals.Text alone has inherent limitations in expression, and humans naturally tend to convey information through a combination of multiple modalities. As illustrated in Figure 1, a user might ask, “Did this athlete achieve the highest score?” while providing an accompanying image. Our benchmark systematically incorporates diverse multi-modal integration patterns, designed to encompass a broader spectrum of multi-modal scenarios and better reflect real-world application complexity. 3) Diverse domains (sports, movies, cartoon, egocentric videos and anomaly detection) ensuring ecological validity. This structure captures the spectrum of real-world retrieval intents.
(3) High Reliability via Human-Centric Design. Unlike synthetic benchmarks, our tasks are meticulously curated through human annotation of 1800 query samples. This ensures query-moment relevance and temporal boundary accuracy, addressing the reliability crisis in automated benchmark generation.
To assess the performance of the current video retrievers, we evaluate three lines of different baselines: clip-style encoders, MLLM-based encoders, and composed video encoders. Our experiments reveal that while these models perform well in specific domains, they lack robustness when tackling more complex, real-world tasks. To address this limitation, we propose a robust embedding framework, V-Embedder, which leverages MLLM Wang et al. (2025) as its backbone and introduces a synthetic data generation framework to train a more generalizable multi-modal video retriever. Our results show that the proposed model delivers robust performance with minimal training data across a wide range of retrieval tasks, highlighting its capability to effectively handle complex multi-modal queries.
In summary, our contributions are threefold:
1. We introduce MomentSeeker, a comprehensive benchmark to evaluate retrieval models’ performance in handling general LVMR tasks.
2. We conduct extensive experiments with various popular multimodal retrievers based on our benchmark, whose results highlight the challenges of LVMR and limitations for existing methods.
3. We further provide an MLLM-based LVMR retriever, V-Embedder, fine-tuned on synthetic data, which demonstrates strong performance on our benchmark.
Figure 2: Overview of the MomentSeeker benchmark. MomentSeeker comprises four meta-tasks, leading to 18 distinct retrieval tasks and 1.8K queries. Each meta-task is illustrated with one or two examples.
2 Related Works
2.1 Video Understanding Benchmarks
As shown in Table 1, current video retrieval benchmarks include MSR-VTT Xu et al. (2016), MSVD Chen and Dolan (2011), VATEX Wang et al. (2020), and ActivityNet Heilbron et al. (2015). These benchmarks evaluate cross-modal retrieval by using video captions as queries to rank videos within a candidate pool. Temporal grounding benchmarks like QVHighlights Lei et al. (2021), Charades-STA Gao et al. (2017), THUMOS14 Jiang et al. (2014), and DiDemo Hendricks et al. (2017) attempt to address moment retrieval but are limited to short videos (under three minutes). This leaves a gap in handling long-form content, where events may require multi-step reasoning.
There are also long-video understanding benchmarks, such as Video-MME Fu et al. (2024) and MLVU Zhou et al. (2025). However, these benchmarks focus on video question answering and do not directly evaluate the representational capability of video encoders. In contrast, our proposed MomentSeeker benchmark is designed for long-video moment retrieval, assessing a model’s retrieval ability within a single long video and addressing the limitations of existing benchmarks.
2.2 Multi-modal Embedding Models
Most video retrievers Li et al. (2022b); Wang et al. (2022a); Zhu et al. (2023) use a dual-tower architecture like CLIP Radford et al. (2021) and train with contrastive learning on large datasets. InternVideo Wang et al. (2022b) integrates generative and discriminative self-supervised learning via masked video modeling and contrastive learning. InternVideo2 Wang et al. (2025) expands upon this by incorporating larger models, featuring an encoder with up to 6 billion parameters. LanguageBind Zhu et al. (2023) aligns multiple modalities in a shared language space, while VAST Chen et al. (2023) integrates vision, audio, and text encoders to have a thorough understanding. These models capture spatiotemporal information, forming the basis of video understanding.
In image retrieval field, several studies Zhang et al. (2024); Zhou et al. (2024); Wei et al. (2025) have explored multi-modal input for composed image retrieval. For example, MagicLens Zhang et al. (2024) uses open-ended instructions to capture complex semantic relationships and trains a composed image retriever, while MM-Ret Zhou et al. (2024) explores further with more effective training data. Inspired by this, CoVR Ventura et al. (2024) applies the same pipeline in video retrieval, generating video-text-video triplets from paired videos with similar captions, and provides a BLIP-style Li et al. (2022a) composed video retriever. These advancements expand the capabilities of foundation models in multi-modal understanding, broadening task adaptability.
Furthermore, with the rise of large language models, some works attempt to transform multi-modal large models (MLLMs) into universal embedders to construct more powerful representations, such as E5-V Jiang et al. (2024) and MLLM2Vec Jiang et al. (2025). However, in the video domain, a robust universal embedder has yet to emerge. For this reason, we propose V-Embedder, which outperforms existing video embedders across multiple tasks through synthetic-data-driven training.
3 Benchmark: MomentSeeker
3.1 Overview
We present MomentSeeker, a comprehensive benchmark to evaluate retrieval models’ performance in handling general long-video moment retrieval (LVMR) tasks. Our framework systematically evaluates model capabilities through three dimensions: 1). Question Types: Descriptive captions vs. natural language queries. 2). Query Modalities: Text-only, text+image, or text+video inputs to support real-world multi-modal search. 3). Data Domains: Five diverse domains (sports, egocentric, movies, cartoons, anomalies) to test generalization. These axes define 18 retrieval scenarios, consolidated into four meta-tasks that probe distinct capabilities:
  • Caption Alignment: Cross-modal matching between caption and video moments.
  • Moment Search: Text-only localization according to instructional queries.
  • Image-Conditioned Moment Search: Retrieval guided by text and image exemplars.
  • Video-Conditioned Moment Search: Retrieval guided by text and video content.
We formulate the moment retrieval task in a long-video-friendly manner. For each video, we extract the ground truth clips for all questions, while the remaining video segments are randomly cropped (2 to 30s) as candidate clips. The embedding models compress the query into a vector and the candidate clips into a set of vectors. The candidate with the highest dot product is selected as the prediction for evaluation. We measure Recall@1 and mAP@5 to reflect the percentage of top candidates matching the ground truth. Figure 2 provides an overview and examples of MomentSeeker.
[…]
Table 1: Comparison with existing video retrieval benchmarks (top) and moment retrieval benchmarks (bottom). INS=Support Instructional Queries, MR=Moment Retrieval, RET=Retrieval, Q=Query, MOD=Modality, Len=Length, VID=Video, Dur=Duration, GT=Ground Truth.
Benchmark Domain INS MR RET Q MOD Avg. Q Len. #VID/ #Clip Min/Max VID Dur. (s) Sum VID Dur. (h) Avg. GT Dur. (s)
MSR-VTT Open $\checkmark$ T 9.42 1000/1000 $9.97$/$30.0$ $4.2$ $15.1$
MSVD Open $\checkmark$ T 7.27 670/670 $1.9$/$60.0$ $1.9$ $10.0$
VATEX Open $\checkmark$ T 14.52 6000/6000 $1.3$/$10.1$ $2.3$ $9.7$
DiDeMo Flickr $\checkmark$ T 29.42 1036/1036 $21.96$/$180.0$ $13.8$ $49.8$
CharadesSTA Activity $\checkmark$ $\checkmark$ T 6.32 1334/3720 $7.6$/$72.1$ $8.2$ $6.9$
QVHighlights Vlog/News $\checkmark$ $\checkmark$ T 10.45 476/1529 $124.0$/$150.0$ $41.6$ $7.9$
THUMOS14 Action $\checkmark$ $\checkmark$ T 1.00 216/3460 $15.0$/$1673.4$ $12.8$ $5.9$
MomentSeeker Open $\checkmark$ $\checkmark$ $\checkmark$ T/T+V/T+I 13.24 268/9976 $5.0$/$7108.9$ $39.9$ $14.9$
3.2 Task Definition
To achieve video content retrieval, we first have human experts annotate precise answer intervals within each video, which serve as ground truth. The remaining segments are randomly split into short candidate clips, forming a retrieval pool. The goal is to retrieve the most relevant clip from a pool of candidates based on a given query. Formally, given a query $q$ and a set of N candidate clips $C=\{c_{\text{1}},c_{\text{2}},\dots,c_{\text{N}}\}$, the retrieval process is defined as:
$c^{*}=\arg\max_{c_{\text{i}}\in C}S(q,c_{\text{i}}),$
where $S(q,c_{\text{i}})$ denotes the similarity score between the query $q$ and candidate clip $c_{\text{i}}$.
In the domain of video retrieval, tasks are categorized into four meta-tasks: Moment Search (MS), Caption Alignment (CA), Image-conditioned Moment Search (IMS), and Video-conditioned Moment Search (VMS). Each of these tasks presents unique challenges and serves distinct purposes in the broader context of multimedia understanding and retrieval. Below, we elaborate on each task.
3.2.1 Caption Alignment
Caption Alignment is defined as retrieving the most relevant segment from a set of candidate videos based on a textual video description $q_{\text{caption}}$ Caption alignment is a fundamental task in video retrieval, aiming to match a textual description with the corresponding video segment. This task requires the system to bridge the gap between explicit textual and visual modalities, serving as the foundation of video retrieval.
3.2.2 Moment Search
Moment Search is defined as retrieving the most relevant segment from a set of candidate videos based on a instructional natural language query $q_{\text{inst}}$ Unlike the straightforward matching in caption alignment, Moment search requires the system to comprehend the natural language instructions, which can be semantically ambiguous or context-dependent. The system must map these instructions to specific visual content within a video, recognizing objects and actions while also comprehending the deeper semantics of the temporal sequence of events.
3.2.3 Image-conditioned Moment Search
Image-conditioned Moment Search is defined as retrieving the most relevant segment from a set of candidate videos based on a natural language query $q_{\text{inst}}$ combined with an image $I$, denoted as $q=\{q_{\text{inst}},I\}$. Image-conditioned moment search introduces an additional layer of complexity by incorporating both textual and visual information in the query. The system must not only understand the natural language query but also interpret the visual context provided by the image. This task requires the integration of multi-modal information, where the image may provide contextual clues that are not explicitly mentioned in the text.
3.2.4 Video-conditioned Moment Search
Video-conditioned Moment Search is defined as retrieving the most relevant segment from a set of candidate videos based on a natural language query $q_{\text{inst}}$ combined with a reference video V, denoted as $q=\{q_{\text{inst}},V\}$. Video-conditioned moment search is perhaps the most complex of the four tasks, as it requires the system to process and understand both a textual query and a reference video. The reference video provides a rich source of contextual information, but it also introduces challenges related to temporal alignment and the extraction of relevant features from the video. The challenge lies in effectively leveraging the temporal and visual information from the reference video to enhance the retrieval process.
Each of these meta-tasks presents unique challenges that push the boundaries of current video retrieval systems. By addressing these challenges, retrievers will benefit more to downstream tasks.
4 V-Embedder: A Flexible and Powerful Video Embedder
Given the diverse range of tasks outlined in the previous section, existing state-of-the-art embedding models Wang et al. (2025); Zhu et al. (2023); Zhou et al. (2024) struggle to offer comprehensive support. To address this, we introduce V-Embedder, a MLLM-based video embedding framework that leverages synthetic data for task-flexible, multi-modal video retrieval. Unlike traditional clip-style retrievers Wang et al. (2025); Zhu et al. (2023), which process textual and visual inputs separately, V-Embedder utilizes a MLLM as its embedding backbone and dynamically integrates textual descriptions with conditioned image frames based on specific task instructions. This unique approach enables fixed-dimensional embeddings optimized for diverse applications, from caption alignment to visual-conditioned moment search. The overall training pipeline consists of two key components: synthetic data generation and contrastive learning.
Synthetic Data Generation.
Due to the scarcity of long-video datasets with transition-specific textual annotations, we use a retrieval-based approach to mine similar video pairs from a large candidate pool. A VLM Wang et al. (2024a) generates transformation instructions describing transitions between video pairs.
To ensure diverse training data, we integrate text-text and video-video retrieval methods, following prior works Zhou et al. (2024); Wang et al. (2024e). For each video $v_{\text{i}}$ and caption $t_{\text{i}}$, we compute video and text embeddings $\mathbf{e}_{v}^{(i)}$ and $\mathbf{e}_{t}^{(i)}$, selecting the top-$k$ ranked pairs from both methods to create a diverse dataset: $\mathcal{D}=\{(V_{1},V_{2}),(V_{1},V_{3}),\dots\}$. A powerful pre-trained VLM Wang et al. (2024a), helps distinguish globally similar yet locally varying videos. A LLM Team (2024) generates relational descriptions, creating flexible, multi-modal training triplets: $\mathcal{D}=\{(V_{1},T_{1},V_{2}),(V_{1},T_{2},V_{3}),\dots\}$.
This approach generates multi-modal data, enhancing the model’s flexibility and generalization for downstream tasks.
Contrastive Training Framework.
[…]
Within this framework, each multi-modal query embedding $\mathbf{e}_{\text{i}}$ is paired with a positive candidate embedding $\mathbf{e}_{\text{i}}^{+}$. The contrastive loss is:
$\mathcal{L}=-\frac{1}{N}\sum_{\text{i}=1}^{N}\log\frac{\exp\left(\mathrm{sim}(\mathbf{e}_{\text{i}},\mathbf{e}_{\text{i}}^{+})/\tau\right)}{\sum_{\text{j}=1}^{N}\exp\left(\mathrm{sim}(\mathbf{e}_{\text{i}},\mathbf{e}_{\text{j}})/\tau\right)}$
To increase the number of in-batch random negatives and enhance the performance of the embedding model, we integrate GradCache Jiang et al. (2025); Gao et al. (2021), which splits inputs into small chunks and optimizes backpropagation.
5 Experiments
5.1 Annotation Overflow
Table 2: Results on the MomentSeeker benchmark. We report R@1 scores averaged across meta-tasks, with all baselines evaluated in a zero-shot setting. Additionally, we provide mAP@5 scores and present detailed performance breakdowns for both metrics on individual sub-tasks in Appendix 8.1.
Method Backbone # Params Meta-Task Average Score Overall
CA MS IMS VMS
number of tasks - - 5 5 4 4 18
CLIP-style Video Encoder
MM-Ret Zhou et al. (2024) CLIP-Base 149M 23.2 15.4 10.5 10.5 14.9
MagicLens Zhang et al. (2024) CLIP-Large 428M 9.0 2.4 3.2 2.8 4.4
UniIR Wei et al. (2025) CLIP-Large 428M 25.0 15.2 6.4 0.0 10.9
LanguageBind Zhu et al. (2023) CLIP-Large 428M 39.6 16.4 3.2 0.0 14.8
InternVideo2 Wang et al. (2025) ViT 1B 44.6 18.2 4.8 0.0 16.9
Composed Video Encoder
CoVR Ventura et al. (2024) BLIP-Large 588M 25.8 17.4 12.3 12.3 17.1
MLLM-based Embedding Model
MLLM2VEC Jiang et al. (2025) LLaVA-1.6 8.4B 6.4 6.2 3.0 3.0 4.7
E5V Jiang et al. (2024) LLaVA-1.6 8.4B 25.8 16.8 6.2 5.2 13.5
V-Embedder InternVideo2-Chat 8B 42.2 20.4 15.0 15.8 23.3
We recruit graduate students in computer science to annotate the dataset. Annotators watch videos and perform: (1) Moment Search (MS): posing a question and marking relevant segments, (2) Image-Conditioned Moment Search (IMS): selecting a frame, posing a question, and marking relevant segments. To balance quality and efficiency, MS and IMS are manually labeled, while Video-Conditioned Moment Search (VMS) is auto-generated by extending IMS frames with a random temporal window. Annotation guidelines are in Appendix 8. Qwen2VL-72B Wang et al. (2024a) generates CA captions for ground truth clips. Annotators cross-examined the dataset to correct errors.
5.2 Experiment Settings
5.2.1 Baselines
Our baselines are grouped into three categories:
CLIP-style Video Encoders. This category includes methods such as MagicLens Zhang et al. (2024), MegaPairs Zhou et al. (2024), UniIR Wei et al. (2025), InternVideo2 Wang et al. (2025), and LanguageBind Zhu et al. (2023). These approaches leverage joint training of video and text to efficiently extract and align cross-modal features.
Composed Video Encoder Approaches. Represented by COVR Ventura et al. (2024), these methods simultaneously consider both text and video queries to search for relevant videos in a database, enabling a more integrated query handling.
MLLM-based Multi-Modality Encoders. Methods like E5V Jiang et al. (2024) and MLLM2Vec Jiang et al. (2025) fall into this category. They utilize large-scale pre-trained multi-modal models to handle flexible modality compositions (e.g., video + text or text alone) and generating multi-modal embeddings.
5.2.2 Implementation Details
Evaluation.
For each video, we extract the ground truth video clips of all questions, fully retaining them. The left parts of the video is randomly cropped (2 to 30s) as the candidate video clips. Each query’s candidate clips are sourced exclusively from its corresponding video.
Following previous works Jiang et al. (2025, 2024), we incorporated instructions (e.g., “Represent the given question in one word.” and “Represent the video in one word.”) to guide the model in generating informative embeddings for queries and candidates. All baselines are reproduced using their official implementations with default settings. Instruction usage remains consistent with each baseline’s original configuration. For image-based baselines, we use the middle frame of the video as input. CoVR Ventura et al. (2024) follows its original configuration with 15 frames at a resolution of 384, while LanguageBind, InternVideo2 Wang et al. (2025), and V-Embedder each use 8 frames at a resolution of 224.
V-Embedder Training.
For synthetic data, we use LanguageBind Zhu et al. (2023) encoders to mine training samples from the WebVid Bain et al. (2022) subset of LVD Xiong et al. (2024). We employ both text-text and video-video retrievers to retrieve the top-5 results each, which are then merged. Transition text labeling is performed by Qwen2VL-72B Wang et al. (2024a) as the MLLM annotator and LLaMA3-8B Team (2024) as the LLM query generator.
For training, we initialize our model from the InternVideo2-Chat(8B) Wang et al. (2025) checkpoint and fine-tune it using LoRA with a rank of 32. After applying GradCache Gao et al. (2021), we train with a batch size of 480 and a learning rate of 5e-4 for 80 steps. The video input consists of 8 frames at a resolution of 224, while images are treated as static videos with the same dimensions.
5.2.3 Evaluation Metrics
We evaluated performance using Recall@1 and MAP@5, with each query ranking video clips from the same long video. Recall is based on exact matches, while MAP considers Intersection over Union (IoU) with ground truth to account for partial correctness, ensuring a balanced evaluation.
5.3 Main Results
The overall evaluation results for all baselines in the MomentSeeker benchmark are shown in Table 2. Individual performances are reported for each task, while average score of Recall@1 are provided. From the results, we derive five primary conclusions:
1. CLIP-style video encoders perform well on text-based retrieval tasks but struggle with multi-modal retrieval. InternVideo2 achieves an R@1 of 44.6 on CA and 18.2 on MS but drops to 4.8 on IMS and 0.0 on VMS. Similarly, LanguageBind, the second-best baseline on CA and MS, achieves 39.6 on CA and 16.4 on MS but also falls to 3.2 on IMS 0.0 on VMS. This suggests that despite their strong temporal modeling, these models fail to establish effective multi-modal combination and alignment.
2. Composed Video Retriever excels in multi-modal tasks but struggles with text-based retrieval robustness. CoVR achieves a score of 12.3 on the VMS benchmark but lags significantly in CA performance, scoring 25.8 compared to InternVideo2’s 44.6. This highlights its suboptimal balance between compositional reasoning and general retrieval capabilities.
3. Composed image retrievers demonstrate unexpected generalization to video tasks. MM-Ret scores 10.5 on both IMS and VMS benchmarks, outperforming specialized video encoders like LanguageBind, which scores only 3.2 on IMS. However, its reliance on single-frame inputs restricts temporal understanding, resulting in an overall subpar performance of 17.1.
4. MLLM-based embedding models dominate instruction-based tasks. The MLLM-powered E5V model surpasses video encoder LanguageBind on MS and VMS tasks, achieving 16.8 and 5.2 respectively against LanguageBind’s 16.4 and 0.0. However, E5V trails substantially in CA performance with 22.3 versus LanguageBind’s 39.6, underscoring persistent challenges in fine-grained temporal alignment.
5. V-Embedder achieves top performance across all tasks, demonstrating strong generalization in both traditional retrieval (CA) and complex instructional scenarios (MS, IMS, VMS). Compared to video foundation models, our model performs comparably to InternVideo2 on the traditional CA task. In tasks requiring complex instruction understanding, our model significantly surpasses these baselines, outperforming the strongest existing baseline by 2.2 points on MS and by over 10 points on IMS and VMS, leading to an overall 6.4 points advantage over InternVideo2. This highlights that our MLLM-Embedder effectively balances temporal modeling, cross-modal alignment, and compositional reasoning, making it a strong and general model.
[…]
Model MSR-VTT VATEX
CLIP Radford et al. (2021) 30.4 -
ViCLIP Wang et al. (2024d) 42.4 -
InternVideo-L Wang et al. (2022a) 40.7 49.5
UMT-L Li et al. (2024) 40.7 -
VideoCoCa-g Yan et al. (2023) 34.4 53.2
VideoPrism-g Zhao et al. (2024) 39.7 62.5
[…]
V-Embedder 43.4 67.7
5.4 Dataset Cross-validation
To assess the generalization capability of V-Embedder, we evaluate its zero-shot performance on traditional video retrieval benchmarks: MSRVTT Xu et al. (2016) and VATEX Wang et al. (2020). As shown in Table 3, V-Embedder achieves state-of-the-art performance on VATEX with a score of 67.7, significantly outperforming VideoPrism-g by around 5 points. While slightly trailing LanguageBind on MSR-VTT, it still surpasses all other baselines. These results highlights V-Embedder’s robust performance across conventional datasets.
5.5 Impact of Training Data
To assess training data impact, we compare our generated data with InternVid Wang et al. (2024d) under equal training steps (Table 5). Experiments demonstrate that the generated data is more efficient than the text-video dataset. This improvement is primarily attributed to its instruction-friendly design and multi-modal support, making it well-suited for handling complex tasks.
5.6 Impact of Retrievers
Table 4: Performance comparison of different retrieval strategies. “Hybrid” refers to the combination of Text-Text and Video-Video retrievers.
Retriever CA MS IMS VMS Overall
Hybrid 42.2 33.5 15.0 15.8 26.6 -
Text-Text 41.9 33.2 14.2 15.1 26.1 -1.8%
Video-Video 41.3 32.9 14.0 14.8 25.8 -3.0%
Table 5: Performance comparison of InternVid Wang et al. (2024d) and our synthetic data.
Data CA MS IMS VMS Overall
InternVid 28.8 18.8 9.5 6.5 15.9
Synthetic 42.2 20.4 15.0 15.8 23.3
To evaluates the impact of different retrieval strategies on model performance, we compare text-text retriever, video-video retriever and our hybrid retriever. As shown in Table 4, using single-modality retrievers yields suboptimal results: the text-text retriever achieves 26.1 overall score while the video-to-video variant performs 25.8 overall score. The hybrid approach combining both retrievers demonstrates superior performance (26.6 AVG), suggesting that cross-modal retrieval enriches training data diversity, whcih leads to more robust video representations.
6 Conclusion
In this paper, we introduce MomentSeeker, a benchmark for evaluating retrieval models in LVMR. MomentSeeker offers three main advantages: it focuses on long videos, covers a wide range of task categories and application scenarios, and ensures reliability through human annotation. We also demonstrates strong performance using a fine-tuned MLLM-based LVMR retriever. Extensive experiments highlight the challenges of LVMR and the limitations of existing methods, with resources shared to advance future research.
7 Limitations
This paper focuses on moment retrieval within long videos and presents a comprehensive benchmark evaluation. For the model, we employ synthetic data to train a MLLM capable of understanding complex instructions and integrating multiple modalities. However, compared to a standalone encoder Wang et al. (2025); Zhu et al. (2023), the MLLM has more parameters, resulting in longer inference time. In the future, we plan to explore techniques such as model distillation to reduce model size and inference cost.
Ethics Statement
Our benchmark video data is sourced from publicly available datasets  Mangalam et al. (2023); Zhou et al. (2025); Wang et al. (2024b); Sultani et al. (2019), which have undergone rigorous screening by the original teams to remove harmful content. Despite our best efforts, we acknowledge that these videos may not be entirely comprehensive or free from omissions. Furthermore, we strongly discourage the use of V-Embedder models for encoding or retrieving sensitive content.

### v3

@@
 Retrieval augmented generation (RAG) holds great promise in addressing challenges associated with long video understanding.
-These methods retrieve useful moments from long videos for their presented tasks, thereby enabling multimodal large language models (MLLMs) to generate high-quality answers in a cost-effective way.
+These methods retrieve useful moments from long videos for their presented tasks, thereby enabling multi-modal large language models (MLLMs) to generate high-quality answers in a cost-effective way.
 In this work, we present MomentSeeker, a comprehensive benchmark to evaluate retrieval models’ performance in handling general long-video moment retrieval (LVMR) tasks.
@@
 We further fine-tune an MLLM-based LVMR retriever on synthetic data, which demonstrates strong performance on our benchmark.
-We perform extensive experiments with various popular multimodal retrievers based on our benchmark, whose results highlight the challenges of LVMR and limitations for existing methods.
+We perform extensive experiments with various popular multi-modal retrievers based on our benchmark, whose results highlight the challenges of LVMR and limitations for existing methods.
 Our created resources is available at https://github.com/yhy-2000/MomentSeeker.
@@
 Given a natural language query $\textbf{q}_{\textbf{T}}$ (optionally accompanied by an image $\textbf{q}_{\textbf{I}}$ or video $\textbf{q}_{\textbf{V}}$), models should search the provided video to precisely identify the most relevant clips.
-We marked the parts that directly answer the query with red circles.
+The ground truth clip is pointed by arrows, and we marked the parts that directly answer the query with red circles.
 Long video understanding remains a significant challenge Fu et al.
@@
 1. We introduce MomentSeeker, a comprehensive benchmark to evaluate retrieval models’ performance in handling general LVMR tasks.
-2. We conduct extensive experiments with various popular multimodal retrievers based on our benchmark, whose results highlight the challenges of LVMR and limitations for existing methods.
+2. We conduct extensive experiments with various popular multi-modal retrievers based on our benchmark, whose results highlight the challenges of LVMR and limitations for existing methods.
 3. We further provide an MLLM-based LVMR retriever, V-Embedder, fine-tuned on synthetic data, which demonstrates strong performance on our benchmark.
@@
 Furthermore, with the rise of large language models, some works attempt to transform multi-modal large models (MLLMs) into universal embedders to construct more powerful representations, such as E5-V Jiang et al.
-(2024) and MLLM2Vec Jiang et al.
+(2024) and VLM2VEC Jiang et al.
 (2025).
@@
 We report R@1 scores averaged across meta-tasks, with all baselines evaluated in a zero-shot setting.
-Additionally, we provide mAP@5 scores and present detailed performance breakdowns for both metrics on individual sub-tasks in Appendix 8.1.
-Method Backbone # Params Meta-Task Average Score Overall
+For image-based models (MM-Ret, MagicLens, UniIR, VLM2VEC, E5V), we use the middle frame of videos as input; other models follow their original optimal settings.
+Additionally, we provide mAP@5 scores and detailed performance on individual sub-tasks in Appendix 8.1.
+Method Backbone # Params Frames Meta-Task Average Score Overall
 CA MS IMS VMS
-number of tasks - - 5 5 4 4 18
-CLIP-style Video Encoder
-MM-Ret Zhou et al. (2024) CLIP-Base 149M 23.2 15.4 10.5 10.5 14.9
-MagicLens Zhang et al. (2024) CLIP-Large 428M 9.0 2.4 3.2 2.8 4.4
-UniIR Wei et al. (2025) CLIP-Large 428M 25.0 15.2 6.4 0.0 10.9
-LanguageBind Zhu et al. (2023) CLIP-Large 428M 39.6 16.4 3.2 0.0 14.8
-InternVideo2 Wang et al. (2025) ViT 1B 44.6 18.2 4.8 0.0 16.9
+number of tasks - - - 5 5 4 4 18
+CLIP-style Visual Encoder
+MM-Ret Zhou et al. (2024) CLIP-Base 149M 1 23.2 15.4 10.5 10.5 14.9
+MagicLens Zhang et al. (2024) CLIP-Large 428M 1 9.0 2.4 3.2 2.8 4.4
+UniIR Wei et al. (2025) CLIP-Large 428M 1 25.0 15.2 6.4 0.0 10.9
+LanguageBind Zhu et al. (2023) CLIP-Large 428M 8 39.6 16.4 3.2 0.0 14.8
+InternVideo2 Wang et al. (2025) ViT 1B 8 44.6 18.2 4.8 0.0 16.9
 Composed Video Encoder
-CoVR Ventura et al. (2024) BLIP-Large 588M 25.8 17.4 12.3 12.3 17.1
+CoVR Ventura et al. (2024) BLIP-Large 588M 15 25.8 17.4 12.3 12.3 17.1
 MLLM-based Embedding Model
-MLLM2VEC Jiang et al. (2025) LLaVA-1.6 8.4B 6.4 6.2 3.0 3.0 4.7
-E5V Jiang et al. (2024) LLaVA-1.6 8.4B 25.8 16.8 6.2 5.2 13.5
-V-Embedder InternVideo2-Chat 8B 42.2 20.4 15.0 15.8 23.3
+VLM2VEC Jiang et al. (2025) LLaVA-1.6 8.4B 8 6.4 6.2 3.0 3.0 4.7
+E5V Jiang et al. (2024) LLaVA-1.6 8.4B 8 25.8 16.8 6.2 5.2 13.5
+V-Embedder InternVideo2-Chat 8B 8 42.2 20.4 15.0 15.8 23.3
 We recruit graduate students in computer science to annotate the dataset.
@@
 Methods like E5V Jiang et al.
-(2024) and MLLM2Vec Jiang et al.
+(2024) and VLM2VEC Jiang et al.
 (2025) fall into this category.
@@
 All baselines are reproduced using their official implementations with default settings.
+We evaluate the video models (LanguageBind and InternVideo2) using an input of uniformly sampled 8 frames, while COVR follows its default setting of 15 frames.
+For image models (E5V, VLM2Vec, UniIR and MagicLens), we use the temporally middle frame as the video input.
 Instruction usage remains consistent with each baseline’s original configuration.
@@
 1.
-CLIP-style video encoders perform well on text-based retrieval tasks but struggle with multi-modal retrieval.
-InternVideo2 achieves an R@1 of 44.6 on CA and 18.2 on MS but drops to 4.8 on IMS and 0.0 on VMS.
-Similarly, LanguageBind, the second-best baseline on CA and MS, achieves 39.6 on CA and 16.4 on MS but also falls to 3.2 on IMS 0.0 on VMS.
-This suggests that despite their strong temporal modeling, these models fail to establish effective multi-modal combination and alignment.
-2.
-Composed Video Retriever excels in multi-modal tasks but struggles with text-based retrieval robustness.
-CoVR achieves a score of 12.3 on the VMS benchmark but lags significantly in CA performance, scoring 25.8 compared to InternVideo2’s 44.6.
-This highlights its suboptimal balance between compositional reasoning and general retrieval capabilities.
+All baseline models exhibit suboptimal performance on our benchmark, underscoring the challenge of the tasks.
+Even state-of-the-art video retrievers like InternVideo2, which achieves 57.9 Recall@1 on Didemo, show drastically reduced effectiveness for long-video temporal grounding (16.9 mean score).
+This stems from two critical challenges: (1) Extended videos contain numerous hard negative segments requiring fine-grained visual-semantic discrimination, and (2) Instructional queries demand deeper temporal reasoning compared to descriptive queries in standard benchmarks.
+Current video-text alignment methods prove inadequate for these human-centric instructional scenarios.
+2. SOTA Models Exhibit Divergent Strengths and Weaknesses Across Tasks.
+  • CLIP-Style Encoders excel in text-based retrieval but falter in multi-modal reasoning.
+For instance, InternVideo2 achieves strong cross-modal alignment (44.6 on CA, 18.2 on MS), yet performs poorly on tasks requiring explicit multi-modal interaction, scoring only 4.8 on IMS and 0.0 on VMS.
+This suggests its temporal modeling capabilities are constrained by isolated modality processing.
+  • Composed Video Retrievers (e.g., CoVR) invert this trend.
+Explicit multi-modal fusion enables CoVR to achieve 12.3 VMS (vs.
+0.0 for CLIP-style baselines), but sacrifices text retrieval robustness, scoring 25.8 on CA versus InternVideo2’s 44.6.
+This highlights a trade-off between multi-modal interaction and cross-modal alignment.
+  • Composed Image Retrievers (e.g., MM-Ret) demonstrate unexpected generalization to video tasks.
+Despite relying on single-frame inputs, MM-Ret scores 10.5 on both IMS and VMS, outperforming specialized video encoders like LanguageBind (3.2 IMS).
+However, its lack of temporal understanding limits overall performance.
+  • MLLM-Based Embedding Models (e.g., E5V) dominate instruction-based tasks.
+E5V surpasses LanguageBind on MS (16.8 vs.
+16.4) and VMS (5.2 vs.
+0.0), leveraging instruction-tuning for complex queries.
+Yet, it underperforms in fine-grained temporal alignment (22.3 CA vs.
+LanguageBind’s 39.6) due to the lack of temporal awareness.
 3.
-Composed image retrievers demonstrate unexpected generalization to video tasks.
-MM-Ret scores 10.5 on both IMS and VMS benchmarks, outperforming specialized video encoders like LanguageBind, which scores only 3.2 on IMS.
-However, its reliance on single-frame inputs restricts temporal understanding, resulting in an overall subpar performance of 17.1.
-4.
-MLLM-based embedding models dominate instruction-based tasks.
-The MLLM-powered E5V model surpasses video encoder LanguageBind on MS and VMS tasks, achieving 16.8 and 5.2 respectively against LanguageBind’s 16.4 and 0.0.
-However, E5V trails substantially in CA performance with 22.3 versus LanguageBind’s 39.6, underscoring persistent challenges in fine-grained temporal alignment.
-5.
-V-Embedder achieves top performance across all tasks, demonstrating strong generalization in both traditional retrieval (CA) and complex instructional scenarios (MS, IMS, VMS).
-Compared to video foundation models, our model performs comparably to InternVideo2 on the traditional CA task.
-In tasks requiring complex instruction understanding, our model significantly surpasses these baselines, outperforming the strongest existing baseline by 2.2 points on MS and by over 10 points on IMS and VMS, leading to an overall 6.4 points advantage over InternVideo2.
-This highlights that our MLLM-Embedder effectively balances temporal modeling, cross-modal alignment, and compositional reasoning, making it a strong and general model.
+Our proposed V-Embedder achieves top performance across all tasks and serves as a strong baseline for future work.
+V-Embedder unifies temporal perception, multimodal reasoning, and instruction understanding into a unified multimodal video retriever, outperforming the previous SOTA baseline CoVR by 6.2 points.
+This improvement leverages our backbone model’s robust video understanding capability, inherited from the pretrained video LLM, combined with fine-tuning on our curated multimodal retrieval dataset—together enabling a breakthrough in multimodal comprehension.
 […]